# From Prompt to Pipeline: Large Language Models for Scientific Workflow Development in Bioinformatics

### v1

@@ 3 Study Design
This study investigates the capabilities of modern LLMs in generating scientific workflows for bioinformatics through natural language prompts. We evaluate the performance of three state-of-the-art LLMs, GPT-4o, Gemini 2.5 Flash, and DeepSeek-V3, in translating high-level user instructions into workflows within two widely used SWSs: Galaxy and Nextflow. Our experimental design focuses on three main research questions: (RQ1) How effectively can LLMs support workflow development?, (RQ2) How do their outputs compare in terms of quality? (RQ3) Which prompting strategies yield the best results?. Through this investigation, we aim to assess the technical capabilities of LLMs and identify best practices in prompt formulation that can support novice and expert users in constructing robust, domain-relevant bioinformatics workflows.
We select Galaxy and Nextflow as representative SWSs due to their widespread adoption within the bioinformatics community [4, 7] and complementary design paradigms. Galaxy provides a user-friendly, graphical web interface that enables researchers, particularly those without programming backgrounds, to design and execute workflows through intuitive drag-and-drop functionality. On the other hand, Nextflow adopts a script-based approach, emphasizing high reproducibility, scalability, and seamless integration with distributed and cloud computing environments. Its flexibility and support for modular, version-controlled workflows make it a preferred choice for expert users engaged in large-scale, automated analyses. For model selection, we consider three advanced LLMs with diverse strengths. GPT-4o is chosen for its demonstrated superiority in code generation and reasoning tasks across multiple domains. Gemini 2.5 Flash is included for its efficiency in producing structured outputs and handling multi-step instructions, traits particularly relevant for workflow synthesis. DeepSeek-V3 is selected for its strong performance in open-source benchmarks and its alignment with code-centric modeling, making it a competitive and transparent alternative to proprietary models. We evaluate each model using its default configuration settings, unless explicitly stated otherwise. This decision reflects the perspective of domain scientists, who may not possess expertise in adjusting model parameters. Requiring manual configuration could introduce unnecessary complexity and create additional barriers to practical use in real-world scientific settings.
@@
W5 nf-core/phyloplace: This bioinformatics performs phylogenetic placement with EPA-NG. [84]
@@
In summary, our approach to RQ1 assesses whether LLMs can accurately understand the scientific context, correctly answer foundational and background questions, and construct usable workflows across a spectrum of bioinformatics tasks. This dual focus on workflow comprehension and generation allows us to systematically measure LLM capabilities in supporting domain scientists, particularly those without programming expertise, in developing robust and reproducible workflows. Detailed results are described in Section 4, including prompt-response summaries, workflow validation details, and model-specific performance insights.
@@
To address RQ2, we systematically evaluate the workflows generated by each LLM along three core dimensions that define the practical utility of scientific workflows in bioinformatics: completeness, correctness, and usability. Our objective is not merely to determine whether an LLM can produce a workflow, but to assess how well the generated workflows align with real-world scientific and technical standards. For this purpose, we select a diverse set of bioinformatics workflows drawn from two widely used scientific workflow systems: Galaxy (graphical) and Nextflow (script-based). These workflows span a range of analysis tasks, including RNA-seq, bacterial genome annotation, quality control of NGS data, and phylogenetic placement, ensuring broad coverage across typical bioinformatics use cases. Each task is benchmarked against a curated, community-reviewed, and publicly documented implementation from authoritative sources such as the Galaxy Training Network and nf-core [85, 36]. To simulate realistic usage scenarios, workflows are generated through conversational prompting using three strategies: instruction-only, role-based, and chain-of-thought. Prompts are applied incrementally, starting with instruction-only and escalating to more structured prompting only when necessary, reflecting typical user behavior where minimal input is preferred.
@@
The first two authors, with over five and nine years of experience in scientific workflows and workflow systems, respectively, manually evaluate each generated workflow. Assessments are based on public documentation, workflow specifications, and verified execution examples. In addition to technical accuracy, we examine narrative clarity, such as explanatory comments and the rationale behind tool selection, which enhances usability, particularly for novice users.
@@
Overall, this evaluation framework is designed to go beyond superficial correctness by measuring how closely LLM-generated workflows approximate expert-designed standards in terms of reproducibility, executability, and scientific relevance. The results offer critical insights into the practical viability of using LLMs for real-world workflow development across diverse tasks and user expertise levels.
@@
Understanding the prompting strategy for interacting with LLMs is essential for achieving accurate, relevant, and reliable responses. Existing research [28, 28, 86, 33] has shown that different prompting techniques, such as zero-shot, few-shot, role-based, and chain-of-thought, lead to varying levels of performance across tasks, particularly in domains requiring structured and sequential reasoning like bioinformatics workflows. Without a deliberate prompting strategy, user queries may yield vague, incomplete, or even incorrect workflows, ultimately limiting the practical utility of LLMs and increasing inefficiency. Therefore, it is essential to evaluate which prompting strategies produce the most accurate, complete, and executable workflows. To address the (RQ3), we conduct an empirical investigation into prompting methods grounded in recent literature on prompt engineering and LLM behavior [e.g., [28, 29, 30, 31, 32, 33, 34, 87, 86]]. Our study first surveys the theoretical foundations of these strategies and their prior success in domains involving procedural generation, planning, and code synthesis. We then experimentally evaluate different prompting techniques across diverse workflow scenarios, refining them through iterative testing. In Section 4, we summarize our findings, providing actionable insights and best practices to help workflow developers optimize their prompting approaches for workflow generation. We provide a practical guideline to support workflow developers who may not possess expertise in prompt engineering or fine-tuning large language models. Recognizing that these users often lack formal training in AI or programming, our guideline is designed to be accessible and actionable, enabling them to effectively adopt our prompting approach for workflow development without requiring deep technical knowledge.
@@ 4 Results
In this section, we present the results of our evaluation of LLMs across three core research questions. We assess their ability to understand fundamental workflow concepts, generate complete and correct scientific workflows, and respond effectively to various prompting strategies. Our findings are based on qualitative and execution-based analyses using workflows from 1.
@@ 4.1 RQ1: To what extent can LLMs assist in the development of scientific workflows?
Motivation: Workflows in bioinformatics are essential for organizing, automating, and scaling complex data analyses. However, designing such workflows typically requires a combination of domain knowledge, technical proficiency, and iterative refinement skills that are often distributed across interdisciplinary teams. For many domain scientists, particularly those from non-computational backgrounds, understanding and constructing workflows using systems like Galaxy or Nextflow presents a significant challenge. SWSs involve abstract representations of tasks, parameter configurations, and tools/modules interactions that may be difficult to grasp without prior experience. Recent advancements in LLMs present an opportunity to bridge this gap. LLMs have demonstrated strong capabilities in natural language understanding, code generation, and step-by-step reasoning. Workflows in bioinformatics are essential for organizing, automating, and scaling complex data analyses. However, designing such workflows typically requires a combination of domain knowledge, technical proficiency, and iterative refinement skills that are often distributed across interdisciplinary teams. For many domain scientists, particularly those from non-computational backgrounds, understanding and constructing workflows using systems like Galaxy or Nextflow presents a significant challenge. Recent advancements in LLMs present an opportunity to bridge this gap.
@@
Table 2: Prompts used to evaluate LLMs’ understanding of core concepts in scientific workflows and workflow systems.
@@
In the second phase of our evaluation, we focus on assessing LLMs’ contextual understanding of specific bioinformatics workflows before prompting them to generate complete pipeline implementations. To formulate relevant background questions, we first implement each of the selected workflows by meticulously following the step-by-step instructions provided in the Galaxy Training Network (GTN) and nf-core documentation. These tutorials are typically enriched with biological and computational context, which we carefully analyze to extract the foundational concepts required for informed workflow generation. This preliminary implementation process serves two purposes: (1) it ensures that we, as evaluators, develop a deep understanding of each workflow’s intent and computational logic; and (2) it allows us to derive meaningful, targeted questions that reflect the type of background knowledge a domain scientist would need prior to developing a workflow. The two authors independently review the documentation and execute each workflow, and then collaboratively synthesize key contextual elements into background prompts designed to test the LLMs’ readiness for the task. The complete sets of background questions for workflows based on Galaxy and Nextflow are provided in Table 4 and Table 5, respectively.
@@
In the final phase of our study addressing RQ1, we task the LLMs with generating bioinformatics workflows using Galaxy or Nextflow, depending on the platform associated with each selected use case. This stage is designed to evaluate the extent to which LLMs can synthesize and operationalize workflow knowledge into executable pipelines. Beyond verifying syntactic validity, we focus on whether the generated workflows exhibit appropriate tool selection, logical task sequencing, and adherence to community-established conventions for reproducibility and usability.
Table 6: Prompts used for Galaxy workflow generation across different use cases
@@
We apply this prompting strategy uniformly across all considered workflows and LLMs, ensuring methodological consistency for fair comparison. This design allows us to assess how varying degrees of prompt specificity influence the LLMs’ ability to generate accurate, complete, and usable workflows in both graphical (Galaxy) and script-based (Nextflow) Scientific Workflow Systems. Ultimately, this phase serves as a bridge between conceptual understanding and practical construction, offering insight into the generative capabilities of LLMs within real-world bioinformatics contexts.
Drawing from existing literature on effective prompting strategies (discussed in RQ3), we design and apply structured prompts to guide LLMs in workflow generation. Tables 6 and 7 present representative prompts used to guide the LLMs in generating our workflows. These examples illustrate how varying levels of contextual scaffolding, ranging from instruction-only to role-based and chain-of-thought prompts, are employed to elicit structured and contextually appropriate responses. While the prompts shown are specific to selected case studies, a consistent prompting strategy is applied across all workflows evaluated in this study, as detailed in the results section. The complete set of prompts and corresponding LLM outputs are included in the supplementary material to promote transparency and support reproducibility. All LLMs are queried using default model settings to simulate a realistic usage scenario, where domain experts in bioinformatics may rely on general prompting capabilities without advanced expertise in prompt engineering.
@@
Results of RQ1: To evaluate the LLMs’ foundational understanding of scientific workflows and workflow systems, we present a set of core conceptual prompts (Table 2). All three models, GPT-4o, Gemini 2.5 Flash, and DeepSeek-V3, are able to generate responses that are technically accurate and aligned with the intended meaning of each question. To further assess their effectiveness, we evaluate the responses using six qualitative criteria: correctness, clarity, completeness, terminology, structure, and usability. A comparison of their performance is summarized in Table 4.1. While all models satisfied the correctness and terminology criteria, notable differences emerged across the remaining dimensions. GPT-4o demonstrated strong clarity and accessibility, offering responses that are well-suited for foundational learning and novice users. Gemini 2.5 Flash provided the most comprehensive and academically precise answers, making it ideal for users with intermediate experience or those seeking formal educational material. DeepSeek-V3 delivered contextually accurate responses enriched with practical examples; however, its outputs are occasionally verbose and informally structured, limiting their immediate instructional usability without refinement.
@@
We next evaluate the responses to the platform-specific background prompts designed to assess LLMs’ contextual understanding of Galaxy and Nextflow (Table 3). The evaluation results, summarized in Table 9, reveal distinct strengths and limitations across the three models. GPT-4o produces accurate, well-structured, and beginner-friendly responses characterized by clear language and logical organization. However, it occasionally omits advanced implementation details such as containerization strategies and execution environments, which are important for reproducibility and deployment. Gemini 2.5 Flash delivers the most academically rigorous and comprehensive responses, offering precise explanations that cover tool integration, ecosystem capabilities, and platform-specific components. Its slightly formal tone and structured presentation make it particularly well-suited for intermediate users or instructional contexts. DeepSeek-V3 performs well in technical accuracy and real-world applicability, often referencing specific tools and platforms such as ToolShed and WorkflowHub. However, its responses are frequently verbose, rely heavily on bullet-point formatting, and lack the narrative flow necessary for effective instructional use without additional editing. Overall, Gemini provides the most complete and balanced output, GPT-4o excels in clarity and accessibility for novice users, and DeepSeek-V3 offers valuable practical insight but would benefit from refinement to improve instructional usability.
@@
To evaluate the ability of LLMs to address workflow-specific background questions, we provide a set of foundational biological and technical prompts (Table 4) related to representative Galaxy workflows. The results of this evaluation (Table 10) indicate that Gemini 2.5 Flash consistently provides the most accurate and comprehensive responses. For example, in response to W2–P3 (What is gene annotation, and how is it used to associate peaks with genes?), Gemini offers a detailed explanation encompassing both functional annotation and genomic coordinate mapping, whereas GPT-4o and DeepSeek-V3 provide less insight into the mechanisms linking peaks to nearby genes. Similarly, for W5–P2 (What is Bakta for bacterial genome annotation?), Gemini elaborates on database usage and output formats, while other models only describe its general purpose.
@@
To assess how well LLMs explain workflow-specific background concepts in Nextflow pipelines, we analyze their responses across a curated set of prompts spanning five real-world Nextflow workflows. These prompts covered foundational biological and computational topics relevant to workflows such as nf-core/rnaseq, fetchngs, and phyloplace. Table 11 summarizes the comparative performance across.
Gemini 2.5 Flash addresses complex topics such as metadata retrieval in fetchngs, reference genome handling in phyloplace, and bisulfite alignment nuances in methylseq, often incorporating tool-specific details, flags, and file structure references. For example, Gemini correctly explained the importance of ENA API calls for resolving sample metadata and provided detailed descriptions of output artifacts. Its responses are comprehensive, though sometimes overly verbose or formal, which could be challenging for novice users. GPT-4o, by contrast, prioritizes clarity and accessibility. It provides succinct, beginner-friendly explanations that are logically structured and easy to follow. For example, its explanation of the FASTQ format and the rationale for adapter trimming in RNA-seq workflows is precise and instructional. However, GPT-4o occasionally lacks depth in areas such as the resolution of experiment-level identifiers or the rationale behind multiple HMM profiles in phyloplace. DeepSeek-V3 delivered technically correct responses with a strong practical orientation. It excels in referencing command-line tools, directory structures, and usage examples. In the context of fetchngs, for instance, it correctly described MD5 checksum validation and discussed the role of sample mapping fields. However, its answers often leaned toward being list-heavy and missed some of the conceptual explanations, such as the biological rationale behind bin refinement in MAG analysis or how GAPPA supports phylogenetic visualization. In summary, while all three LLMs perform reasonably well in covering core concepts, Gemini 2.5 Flash is best suited for users seeking comprehensive, GPT-4o excels in readability and instructional clarity, and DeepSeek-V3 offers detailed practical insights but requires refinement for conceptual completeness and consistency.

### v2

@@ 3 Study Design
This study investigates how state-of-the-art LLMs, GPT-4o, Gemini 2.5 Flash, and DeepSeek-V3, can generate bioinformatics workflows from natural language prompts using two widely adopted Scientific Workflow Systems: Galaxy and Nextflow. Our experimental design focuses on three main research questions: (RQ1) How effectively can LLMs support workflow development?, (RQ2) How do their outputs compare in terms of quality? (RQ3) Which prompting strategies yield the best results?. Through this investigation, we aim to assess the technical capabilities of LLMs and identify best practices in prompt formulation that can support novice and expert users in developing robust, domain-relevant bioinformatics workflows. Figure 2 shows the overall approaches of our study methodology.
Figure 2: Overview of our study methodology
We select Galaxy and Nextflow as representative SWSs due to their widespread adoption within the bioinformatics community [4, 7] and complementary design paradigms. For model selection, we consider three advanced LLMs with diverse strengths. GPT-4o is chosen for its demonstrated superiority in code generation and reasoning tasks across multiple domains. Gemini 2.5 Flash is included for its efficiency in producing structured outputs and handling multi-step instructions, traits particularly relevant for workflow synthesis. DeepSeek-V3 is selected for its strong performance in open-source benchmarks and its alignment with code-centric modeling, making it a competitive and transparent alternative to proprietary models. We evaluate each model using its default configuration settings, unless explicitly stated otherwise. This decision reflects the perspective of domain scientists, who may not possess expertise in adjusting model parameters. Requiring manual configuration could introduce unnecessary complexity and create additional barriers to practical use in real-world scientific settings.
@@
W5 nf-core/phyloplace: This workflow performs phylogenetic placement with EPA-NG. [90]
@@
In summary, our approach to RQ1 assesses whether LLMs can accurately understand the scientific context, correctly answer foundational and background questions, and construct usable workflows across a spectrum of bioinformatics tasks. Detailed results are described in Section 4, including prompt-response summaries, workflow validation details, and model-specific performance insights.
@@
To address RQ2, we systematically evaluate the workflows generated by each LLM along three core dimensions that define the practical utility of scientific workflows in bioinformatics: completeness, correctness, and usability. Our objective is not merely to determine whether an LLM can produce a workflow, but to assess how well the generated workflows align with real-world scientific and technical standards. For this purpose, we utilize the workflows listed in Table 1. Each task is benchmarked against a curated, community-reviewed, and publicly documented implementation from authoritative sources such as the GTN and nf-core [80, 36]. To simulate realistic usage scenarios, workflows are generated through conversational prompting using three strategies: instruction-only, role-based, and chain-of-thought. Prompts are applied incrementally, starting with instruction-only and escalating to more structured prompting only when necessary, reflecting typical user behavior where minimal input is preferred.
@@
Two authors, with over five and nine years of experience in scientific workflows and workflow systems, respectively, manually evaluate each generated workflow. Assessments are based on public documentation, workflow specifications, and verified execution examples. In addition to technical accuracy, we examine narrative clarity, such as explanatory comments and the rationale behind tool selection, which enhances usability, particularly for novice users.
@@
Overall, our study is designed to go beyond superficial correctness by measuring how closely LLM-generated workflows approximate expert-designed standards in terms of reproducibility, executability, and scientific relevance. The results offer insights into the practical viability of using LLMs for real-world workflow development across diverse tasks and user expertise levels.
@@
Understanding the prompting strategy for interacting with LLMs is essential for achieving accurate, relevant, and reliable responses. Existing research [28, 91, 33] has shown that different prompting techniques, such as zero-shot, few-shot, role-based, and chain-of-thought, lead to varying levels of performance across tasks. Without a deliberate prompting strategy, user queries may yield vague, incomplete, or even incorrect workflows, ultimately limiting the practical utility of LLMs and increasing inefficiency. Therefore, it is essential to evaluate which prompting strategies produce the most accurate, complete, and executable workflows. To address the (RQ3), we investigate prompting methods grounded in recent literature on prompt engineering and LLM behavior [e.g., [28, 29, 30, 31, 32, 33, 34, 92, 91]]. Our study surveys the theoretical foundations of these strategies and their prior success in domains involving procedural generation, planning, and code synthesis. We then experimentally evaluate different prompting techniques across diverse workflow scenarios, refining them through iterative testing. In Section 4, we summarize our findings, providing actionable insights and best practices to help workflow developers optimize their prompting approaches for workflow generation. We provide a practical guideline to support workflow developers who may not possess expertise in prompt engineering or fine-tuning large language models. Recognizing that these users often lack formal training in AI or programming, our guideline is designed to be accessible and actionable, enabling them to effectively adopt our prompting approach for workflow development without requiring deep technical knowledge.
@@ 4 Results
In this section, we present the results of our evaluation of LLMs across three research questions. We assess their ability to understand fundamental workflow concepts, generate complete and correct scientific workflows, and respond effectively to various prompting strategies. Our findings are based on qualitative and execution-based analyses using workflows from Table 1.
@@ 4.1 RQ1: To what extent can LLMs assist in the development of scientific workflows?
Motivation: Workflows in bioinformatics are essential for organizing, automating, and scaling complex data analyses. However, designing such workflows typically requires a combination of domain knowledge, technical proficiency, and iterative refinement skills that are often distributed across interdisciplinary teams. For many domain scientists, particularly those from non-computational backgrounds, understanding and constructing workflows using systems like Galaxy or Nextflow presents a significant challenge. SWSs involve abstract representations of tasks, parameter configurations, and tools/modules interactions that may be difficult to grasp without prior experience. LLMs have demonstrated strong capabilities in natural language understanding, code generation, and step-by-step reasoning. Recent advancements in LLMs present an opportunity to bridge this gap.
@@
Table 2: Prompts used to evaluate LLMs’ understanding of core concepts in workflows and workflow systems.
@@
In the second phase of our evaluation, we focus on assessing LLMs’ contextual understanding of specific bioinformatics workflows before prompting them to generate complete pipeline implementations. To formulate relevant background questions, we first implement each of the selected workflows by meticulously following the step-by-step instructions provided in the GTN and nf-core documentation. These tutorials are typically enriched with biological and computational context, which we carefully analyze to extract the foundational concepts required for informed workflow generation. This preliminary implementation process serves two purposes: (1) it ensures that we, as evaluators, develop a deep understanding of each workflow’s intent and computational logic; and (2) it allows us to derive meaningful, targeted questions that reflect the type of background knowledge a domain scientist would need prior to developing a workflow. The two authors independently review the documentation and execute each workflow, and then collaboratively synthesize key contextual elements into background prompts designed to test the LLMs’ readiness for the task. The complete sets of background questions for workflows based on Galaxy and Nextflow are provided in Table 4 and Table 5, respectively.
@@
In the final phase of our study addressing RQ1, we task the LLMs with generating workflows using Galaxy or Nextflow, depending on the platform associated with each selected use case. This stage is designed to evaluate the extent to which LLMs can synthesize and operationalize workflow knowledge into executable pipelines. Beyond verifying syntactic validity, we focus on whether the generated workflows exhibit appropriate tool selection, logical task sequencing, and adherence to community-established conventions for reproducibility and usability.
Table 6: Prompts used for Galaxy workflow generation across different use cases.
@@
We apply this prompting strategy uniformly across all considered workflows and LLMs, ensuring methodological consistency for fair comparison. This design allows us to assess how varying degrees of prompt specificity influence the LLMs’ ability to generate accurate, complete, and usable workflows in Galaxy and Nextflow. Ultimately, this phase serves as a bridge between conceptual understanding and practical construction, offering insight into the generative capabilities of LLMs within real-world bioinformatics contexts.
Drawing from existing literature on effective prompting strategies (discussed in RQ3), we design and apply structured prompts to guide LLMs in workflow generation. Tables 6 and 7 present representative prompts used to guide the LLMs in generating our workflows. These examples illustrate how varying levels of contextual scaffolding, ranging from instruction-only to role-based and chain-of-thought prompts, are employed to elicit structured and contextually appropriate responses. While the prompts shown are specific to selected case studies, a consistent prompting strategy is applied across all workflows evaluated in this study, as detailed in the results section. The complete set of prompts and corresponding LLM outputs are included in the supplementary material to promote transparency and support reproducibility.
@@
Results of RQ1:
We present a set of core conceptual prompts to evaluate the LLMs’ foundational understanding of workflows and workflow systems (Table 2). All three models generate technically accurate responses and are aligned with each question’s intended meaning. To further assess their effectiveness, we (two authors, with over five and nine years of experience) evaluate the responses using six qualitative criteria: correctness, clarity, completeness, terminology, structure, and usability. A comparison of their performance is summarized in Table 4.1.
While all models satisfied the correctness and terminology criteria, notable differences emerged across the remaining dimensions. GPT-4o demonstrated strong clarity and accessibility, offering responses that are well-suited for foundational learning and novice users. Gemini 2.5 Flash provided the most comprehensive and academically precise answers, making it ideal for users with intermediate experience or those seeking formal educational material. DeepSeek-V3 delivered contextually accurate responses enriched with practical examples; however, its outputs are occasionally verbose and informally structured, limiting their immediate instructional usability without refinement.
We next evaluate the responses to the platform-specific background prompts designed to assess LLMs’ contextual understanding of Galaxy and Nextflow (Table 3). The evaluation results, summarized in Table 9, reveal distinct strengths and limitations across the three models. GPT-4o produces accurate, well-structured, and beginner-friendly responses characterized by clear language and logical organization. However, it occasionally omits advanced implementation details such as containerization strategies and execution environments, which are important for reproducibility and deployment. Gemini 2.5 Flash delivers the most academically rigorous and comprehensive responses, offering precise explanations that cover tool integration, ecosystem capabilities, and platform-specific components. Its slightly formal tone and structured presentation make it particularly well-suited for intermediate users or instructional contexts. DeepSeek-V3 performs well in technical accuracy and real-world applicability, often referencing specific tools and platforms such as ToolShed and WorkflowHub. However, its responses are frequently verbose, rely heavily on bullet-point formatting, and lack the narrative flow necessary for effective instructional use. Overall, Gemini provides the most complete and balanced output, GPT-4o excels in clarity and accessibility for novice users, and DeepSeek-V3 offers valuable practical insight but would benefit from refinement to improve instructional usability.
@@
To evaluate the ability of LLMs to address workflow-specific background questions, we provide a set of foundational biological and technical prompts (Table 4) related to representative Galaxy workflows.
@@
The results of this evaluation (Table 10) indicate that Gemini 2.5 Flash consistently provides the most accurate and comprehensive responses. For example, in response to W2–P3 (What is gene annotation, and how is it used to associate peaks with genes?), Gemini offers a detailed explanation encompassing both functional annotation and genomic coordinate mapping, whereas GPT-4o and DeepSeek-V3 provide less insight into the mechanisms linking peaks to nearby genes. Similarly, for W5–P2 (What is Bakta for bacterial genome annotation?), Gemini elaborates on database usage and output formats, while other models only describe its general purpose.
@@
To assess how well LLMs explain workflow-specific background concepts in Nextflow pipelines, we analyze their responses using representative Nextflow workflows. These prompts covered foundational biological and computational topics relevant to our workflows. Table 11 summarizes the comparative performance across.
@@
Gemini 2.5 Flash addresses complex topics such as metadata retrieval in fetchngs, reference genome handling in phyloplace, and bisulfite alignment nuances in methylseq, often incorporating tool-specific details, flags, and file structure references. For example, Gemini correctly explained the importance of ENA API calls for resolving sample metadata and provided detailed descriptions of output artifacts. Its responses are comprehensive, though sometimes overly verbose or formal, which could be challenging for novice users. GPT-4o, by contrast, prioritizes clarity and accessibility. It provides succinct, beginner-friendly explanations that are logically structured and easy to follow. For example, its explanation of the FASTQ format and the rationale for adapter trimming in RNA-seq workflows is precise and instructional. However, GPT-4o occasionally lacks depth in areas such as the resolution of experiment-level identifiers or the rationale behind multiple HMM profiles in phyloplace. DeepSeek-V3 delivered technically correct responses with a strong practical orientation. It excels in referencing command-line tools, directory structures, and usage examples. In the context of fetchngs, for instance, it correctly described MD5 checksum validation and discussed the role of sample mapping fields. However, its answers often leaned toward being list-heavy and missed some of the conceptual explanations, such as the biological rationale behind bin refinement in MAG analysis or how GAPPA supports phylogenetic visualization.
In summary, while all three LLMs perform reasonably well in covering core concepts, Gemini 2.5 Flash is best suited for users seeking comprehensive, GPT-4o excels in readability and instructional clarity, and DeepSeek-V3 offers detailed practical insights but requires refinement for conceptual completeness and consistency.